OpenStudy (anonymous):

Can someone please help me with this? Explain how to do it and everything? Which of the following is the slope between the points (1, 5)and (5, 2)? Answer three fourths -negative three fourths four thirds -negative four thirds

OpenStudy (matt101):

The slope is the rise/run. Using your two points, the rise is 5-2=3 units, and the run is 1-5=-4. The slope is therefore -3/4. This means that for every 3 units of rise, there are 4 units of run towards the negative end of the graph.

OpenStudy (anonymous):

Ohhh.... So.. I get it now!! :DD

OpenStudy (anonymous):

But, what about this one? 3--2=5 8-4=4? Which of the following is the slope between the points (3, -2)and (8, 4)? six fifths -negative five sixths -negative six fifths five sixths

OpenStudy (matt101):

You mixed up the math - its the difference between y values and x values between the two points, not between the y and x values of each point (a meaningless number). In this case, you would have: rise=4-(-2)=6 run=8-3=5 slope=rise/run=6/5

OpenStudy (anonymous):

What about these kinds of questions? Identify the correct x intercept and y intercept of the equation 2x - 3y = -18. Answer x intercept = (6, 0); y intercept = (0, -9) x intercept = (9, 0); y intercept = (0, -6) x intercept = (-9, 0); y intercept = (0, 6) x intercept = (-6, 0); y intercept = (0, 9)

OpenStudy (matt101):

The x intercept is the point at which the graph crosses the x axis, which means the y-value of that point must be 0. The y intercept is the point at which the graph crosses the y axis, which means the x-value of that point must be 0. Therefore, to find the x intercept, plug in 0 for y in the equation and solve, and to find the y intercept, plug in 0 for x in the equation and solve. In this case: x intercept (y=0): 2x - 3(0) = -18 2x = -18 x = -9 y intercept (x=0) 2(0) - 3y = -18 -3y = -18 y = 6 The x intercept is (-9,0) and the y intercept is (0,6). Make sense?

OpenStudy (anonymous):

Barely.... I just want to understand 1 thing. Sometimes will the equation till you the value of y or x or will it forever be 0?

OpenStudy (matt101):

It will always be 0.